# I Predict Therefore I Am: Is Next Token Prediction Enough to Learn Human-Interpretable Concepts from Data?

### v1

@@ G.1 Simulation Details
G.2 Experimental Details on LLMs
@@ 27 German $\Rightarrow$ Spanish (Arbeit, trabajo) 22
Appendix H Further Discussion: Observations on LLMs and World Models
H.1 LLMs Mimic the Human World Model, Not the World Itself
@@
H.2 LLMs Versus Pure Vision Models: A Fundamental Difference
@@
Appendix I More Results on Llama-3 and DeepSeek-R1
We conduct additional experiments on recent LLMs, including Llama-3 and DeepSeek-R1, to further evaluate our findings. The experimental setup strictly adheres to the settings described in Section G.2. This enables a comprehensive investigation of our findings, ensuring that the results are thoroughly evaluated and validated across diverse LLM architectures and experimental conditions. Overall, we can see, the product $\mathbf{A}_{s}\times\mathbf{W}_{s}$ approximates the identity matrix, supporting the theoretical findings outlined in Corollary 4.2.

### v2

@@ F.1 Simulation Details
F.2 Experimental Details on LLMs
@@ 27 German $\Rightarrow$ Spanish (Arbeit, trabajo) 22
Appendix G Further Discussion: Observations on LLMs and World Models
G.1 LLMs Mimic the Human World Model, Not the World Itself
@@
G.2 LLMs Versus Pure Vision Models: A Fundamental Difference
@@
Appendix H More Results on Llama-3 and DeepSeek-R1
We conduct additional experiments on recent LLMs, including Llama-3 and DeepSeek-R1, to further evaluate our findings. The experimental setup strictly adheres to the settings described in Section F.2. This enables a comprehensive investigation of our findings, ensuring that the results are thoroughly evaluated and validated across diverse LLM architectures and experimental conditions. Overall, we can see, the product $\mathbf{A}_{s}\times\mathbf{W}_{s}$ approximates the identity matrix, supporting the theoretical findings outlined in Corollary 4.2.